Tiggie's Dog Toys, Inc. reported a debt-to-equity ratio of 1.20 times at the end of 2018. The firm's total assets at year-end were $28.60 million. How much of their assets are financed with debt and how much with equity? (Do not round intermediate calculations. Enter your answer in millions of dollars rounded to 3 decimal places.)
Answer #1

debt-to-equity ratio=debt/equity

Hence debt=1.2*equity

Total assets=debt+equity

28.6=1.2*equity+equity

equity=28.6/(1.2+1)

=$13 million

debt=1.2*equity

=(1.2*13)=$15.6 milion